please help me with this last problem,
Val's mom is 28 years older than she is. In 15 years three times Val's age increased by her mom's age then, will equal 168 years. How old is each now? In 15 years?
Val now-
Mom now-
Val in 15-
Mom in 15-

Answer by Paul(988)

Val's mom = x+28
Val = x
In 15 years Val's age = x+15
In 15 years Val mom's age = x+43
Equation:
3(x+15)+x+43=168
3x+45+x+43=168
4x+88=168
4x=80
x=20

20+28=48

Hence, Val is 20 years old ans Val's mom is 48 years old.
Paul.